3. The Henry’s law constant for CO2 in water at 25ºC is 3.4×10–2 mol/L·atm.
(a) Calculate the concentration of CO2 in a beer can with a partial pressure 3.0 atm over the beer at 25 ºC.

(b) Wascana Lake is located in city of Regina. Given that the partial pressure of CO2 is 3.9 × 10–4 atm, calculate the concentration of CO2 at 25 ºC in Wascana Lake.
1
Expert's answer
2015-12-09T11:57:31-0500
C=k·P(gas)
where
C - the solubility of a gas at a fixed temperature in a particular solvent, mL gas/L;
K - Henry's law constant, mol/L·atm;
Pgas - the partial pressure of the gas, atm.

a) C = 3.0·3.4×10–2=0.102 ml/L
b) C = 3.9·10–4·3.4×10–2=1.3·10-5 ml/L